OpenStudy (anonymous):

for two events, M and N , P(M)=0.4 P(N I M)=0.6, And p(N I M')=0.3 Find P( M' I N) =? anyone?

myininaya (myininaya):

You should have a formula for P(A|B)

myininaya (myininaya):

\[P(A|B)=\frac{P( A and B)}{P(B)}\]

OpenStudy (anonymous):

so how we going to apply this formula to my question?

myininaya (myininaya):

You are ask to find P(M'|N)

myininaya (myininaya):

Look we have: \[P(M' |N)=\frac{P(M' and N)}{P(N)} \text{ and } P(N|M')=\frac{P(N and M')}{P(M')}\] We want to find the first, so we have: \[P(M'|N)=\frac{P(M' and N)}{P(N)}=\frac{P(M')P(N|M')}{P(N)}\]